# Efficient Reinforcement Learning for Optimal Control with Natural Images

### v1

@@ Abstract
Reinforcement learning solves optimal control and sequential decision problems widely found in control systems engineering, robotics, and artificial intelligence. This work investigates optimal control over a sequence of natural images. The problem is formalized, and general conditions are derived for an image to be sufficient for implementing an optimal policy. Reinforcement learning is shown to be efficient only for certain types of image representations. This is demonstrated by developing a reinforcement learning benchmark that scales easily with number of states and length of horizon, and has optimal policies that are easily distinguished from suboptimal policies. Image representations given by overcomplete sparse codes are found to be computationally efficient for optimal control, using fewer computational resources to learn and evaluate optimal policies. For natural images of fixed size, representing each image as an overcomplete sparse code in a linear network is shown to increase network storage capacity by orders of magnitude beyond that possible for any complete code, allowing larger tasks with many more states to be solved. Sparse codes can be generated by devices with low energy requirements and low computational overhead.
Keywords Natural image sequences  $\cdot$ Overcomplete sparse codes  $\cdot$ Optimal policies  $\cdot$ Sufficient statistics
@@ 1 Introduction
Many interesting and complex tasks can be described using natural image sequences. Each image of the sequence corresponds to a particular state of the environment, and choosing an action or applying a control at the current state takes the environment to its next state, and leads to the next image in the sequence. A natural question is then: given an image, what is the best control to apply? Although the answer will depend on the particular task at hand a general framework can be identified. As a simple example, consider moving through a rainforest and taking images of flowers in order to locate a particular species of plant. In this case, each image describes a state of the environment, and the state changes depending on the location of the next image. Each image also has a cost or reward associated with it: an image of the sought-after flower will receive a high reward, while an image without any flowers may receive a low (or even negative) reward. To perform this task well, the choice of control leading to the location of the next image does not necessarily give the best chance of obtaining an image of the sought-after flower at the next time period, rather, it should also lead to a good chance of obtaining such images at future time periods as you move through the rainforest. Planning over a sequence of time periods is the defining characteristic of sequential decision problems and optimal control, and is clearly a requirement of any intelligent system.
@@
Natural images are a high-dimensional dataset with a great deal of expressive power. For some tasks, images may include enough task-dependent information to determine an optimal policy. In this case an image is similar to a belief state in a partially observed Markov decision process, and we will derive the general conditions for an image to be sufficient for implementing an optimal policy. Similar to other types of ecological data, it is well-known that the statistical properties of natural images distinguish them from artificially generated images (Field, 1987, 1994; Ruderman and Bialek, 1994; Ruderman, 1997; Simoncelli and Olshausen, 2001; Hyvärinen et al., 2009; Hosseini et al., 2010). We exploit these statistical properties to find efficient image representations that significantly decrease the computational resources required for finding optimal policies.
The aim of this work is to develop a general framework for the efficient solution of Markov decision processes and optimal control tasks involving natural image sequences. This includes developing a new reinforcement learning benchmark for comparing the performance and efficiency of different image representations. The benchmark must be scalable so the number of states and time periods can easily be increased; and optimal solutions must easily be distinguishable from suboptimal solutions. Scalability of the benchmark is necessary for determining which image representations will perform best over long horizons and for large numbers of states.
@@
Most reinforcement learning is done with synthetic data of limited complexity, and previous work in this area has suggested that more work should be done to extend the paradigm of reinforcement learning to real-world data such as natural images and natural video (Zhang et al., 2018). Past examples of these efforts include using image representations to integrate unsupervised training of deep neural networks into reinforcement learning; specifically for visual navigation (Lange and Riedmiller, 2010), and simulated autonomous driving tasks (Zhang et al., 2021). In order to obtain better policies with neural networks on standard classic control tasks, previous work has also investigated the use of sparse coding and sparse representations in reinforcement learning (Le et al., 2017; Liu et al., 2018; Rafati and Noelle, 2019). Although none of these works considered images or other high-dimensional datasets, and computational efficiency was not investigated. The closest previous work to that presented here is in Loxley (2021), where sparse image representations were shown to significantly decrease the computational resources required for finding optimal policies. However, this work was limited by a small video dataset, which meant that extending results to large numbers of states and long horizons was not possible. Further, only a deterministic optimal control task was considered rather than a general Markov decision process, and the overcomplete sparse representations investigated were very small. In this work, a more general approach to optimal control is taken, and the development of a new reinforcement learning benchmark allows results for image representations to be significantly scaled up by orders of magnitude.
In this work, natural images are represented as overcomplete sparse codes. Statistical dependencies in natural images lead to redundancy, making it possible to obtain an efficient representation using an appropriate encoding scheme such as sparse coding (Daugman, 1988, 1989; Olshausen and Field, 1996, 1997). A sparse code is efficient in the sense of being a low bit-rate (low entropy) representation of a high-dimensional dataset. As shown here, an overcomplete sparse code also turns out to be computationally efficient for finding optimal policies in reinforcement learning. A sparse code for natural images can be generated in many different ways – from simple constructions using multiresolution wavelet bases and autoencoders (Daugman, 1988; Loxley, 2017, 2021), through to heavy-duty approaches that apply deep neural networks (Papyan et al., 2017; Li et al., 2024). The ICA algorithm (Bell and Sejnowski, 1997; Hyvärinen et al., 2009) is another well-know method for generating sparse codes; however, it cannot be applied here as it does not generate overcomplete sparse codes, which are key to this work. Sparse codes may also be generated directly from hardware using purpose built low-power neuromorphic computer chips (Fair et al., 2019).
@@
The structure of this paper is as follows. In Sec. 2, the reinforcement learning methodology is briefly summarized, and the problem of optimal control over a sequence of natural images is formalized. Following this, a new reinforcement learning benchmark is developed step by step. In Sec. 3, results for the reinforcement learning benchmark are presented, and different image representations are compared. A detailed discussion of results and conclusions is given in Sec. 4.
@@ 2 Reinforcement Learning over Natural Image Sequences
Consider a general optimal control task involving a discrete-time dynamical system with a finite number of states and controls (Bertsekas, 2017, 2019). During any time period $k$ the environment is described by a state $i_{k}$ taken from a state-space $S_{k}$, and a control $u_{k}$ is selected from a set of available controls $U_{k}(i_{k})\subset C_{k}$. The state of the environment changes when a controller chooses control $u$ during the current time period, so that state $i$ updates to the successor state $j$ according to the transition probabilities $p_{ij}(u)$ of a controllable Markov chain. The controller can choose different controls depending on the current state of the environment as well as the time period under consideration: the controller’s choice is then described by a policy $\pi=\{\mu_{0},...,\mu_{N-1}\}$, which is a sequence of functions $\mu_{k}$ that map states into controls: $\mu_{k}(i_{k})=u_{k}$, for each state $i_{k}$ at each time period $k$. The cost of being in state $i_{k}$ at time period $k$, choosing control $u_{k}$, then moving to state $i_{k+1}$, is denoted by $g_{k}(i_{k},u_{k},i_{k+1})$. The key quantity of interest is the expected total cost. Given a policy $\pi$ and an initial state $i_{0}$, the expected total cost is the expected value of the sum of costs:
@@
The finite-horizon optimal control task outlined in Eqs. (1) and (2) can be solved exactly using the Dynamic Programming Algorithm, leading to the optimal cost-to-go $J_{k}^{*}$ (this algorithm is called Value Iteration in the case of infinite horizon tasks, and converges to $J^{*}$ for discounted problems with stationary policies). The optimal policy $\mu_{k}^{*}(i)$ can then be found from the following minimization:
@@
by choosing the minimizing control $u^{*}$ in Eq. (3) for each state $i$ and each time period $k$.
Optimal control over natural image sequences has additional issues to address; such as how to represent natural images as states, what information to include in a state to allow for optimal decision-making, and how to include natural images into optimal control algorithms. In the simplest case, each state $i_{k}\in S_{k}$ in an optimal control task becomes a natural image $\phi_{k}(i)\in S_{k}^{\prime}$; where $S_{k}^{\prime}\subset\mathbb{R}^{p}$ so that each image is represented as a $p$-dimensional vector. The cost-to-go can then be described using a parametric function approximator $\bar{J}_{k}(\phi_{k}(i),r_{k})$ that approximates the optimal cost-to-go $J_{k}^{*}$. This can be done by using a neural network architecture that accepts images $\phi_{k}(i)$ as input, and adjusting the network parameters $r_{k}\in\mathbb{R}^{p}$ (called “weights”) to approximate $J_{k}^{*}$ by $\bar{J}_{k}$. When neural network training includes all states at each time period and the neural network is below its memory capacity, $\bar{J}_{k}$ may provide an accurate description of $J_{k}^{*}$. The optimal policy $\mu_{k}^{*}(i)$ can then be found from the following minimization:
$\underset{u\in U_{k}(i)}{\operatorname{min}}\ \sum_{j}p_{ij}(u)\left(g_{k}(i,u,j)+\bar{J}_{k+1}(\phi_{k+1}(j),r_{k+1}^{*})\right),$ (4)
by choosing the minimizing control $u^{*}$ in expression (4) for each state $i$ and each time period $k$.
A key issue to address is what information to include in a state. Any state that includes all necessary information for optimal control is called a sufficient statistic (Bertsekas, 2017). A natural image $\phi_{k}$ is a sufficient statistic, provided $U_{k}(i),p_{ij}(u)$, and $g_{k}(i,u,j)$ in Eq. (4) depend on $i$ or $j$ only through $\phi_{k}(i)$ or $\phi_{k+1}(j)$, respectively. The optimal policy can then be written as
$\mu_{k}^{*}(i)=\bar{\mu}_{k}(\phi_{k}(i)),$ (5)
where $\bar{\mu}_{k}$ is some appropriate function. The most striking consequence of $\phi_{k}$ as a sufficient statistic is the form the transition probabilities now take:
$p_{ij}(u)=\bar{p}(\phi_{k+1}(j)|\phi_{k}(i),u),$ (6)
so that given a current image $\phi_{k}(i)$, selecting control $u$ leads to a new image $\phi_{k+1}(j)$ at the next time period (with some probability). Therefore, the transition probabilities describe an “image generator”. We derive a simple image generator in the next section.
@@
To approximate the optimal cost-to-go $J_{k}^{*}(i)$ using $\bar{J}_{k}(\phi_{k}(i),r_{k})$, the weights $r_{k}$ can be found using Fitted Value Iteration. To start the iteration, we go to the end of the task at $k=N-1$, and make use of the known terminal cost $\bar{J}_{N}(\phi_{N}(i),r_{N})=g_{N}(i)$. The following pair of expressions are then evaluated to yield $\bar{J}_{N-1}(\phi_{N-1}(i),r_{N-1})$:
$\beta_{k}^{i}=\underset{u\in U_{k}(i)}{\operatorname{min}}\ \sum_{j}p_{ij}(u)\left(g_{k}(i,u,j)+\bar{J}_{k+1}(\phi_{k+1}(j),r_{k+1})\right),\ \ \ \ \ \forall i\in S_{k}$ (7)
$\text{minimize}\ \ \ \sum_{i\in S_{k}}(\bar{J}_{k}(\phi_{k}(i),r_{k})-\beta_{k}^{i})^{2},$ (8)
subject $\text{subject to}\ \ \ r_{k}\in\mathbb{R}^{p}.$
@@
This process is continued by iterating backwards from $k=N-2$ to $k=0$, yielding the cost-to-goes $\bar{J}_{N-2},...,\bar{J}_{0}$ required in (4) for finding an optimal policy. The first expression (7) performs the minimization given in Eq. (4) for each of the states using the parametric function $\bar{J}_{k+1}(\phi_{k+1}(j),r_{k+1})$ for the cost-to-go of image $\phi_{k+1}(j)$ at time period $k+1$. This results in the cost-to-go $\beta_{k}^{i}$ for image $\phi_{k}(i)$ at time period $k$. In the second expression (8), the state and cost-to-go pair $(i,\beta_{k}^{i})$ from Eq. (7) is used to fit the parametric function $\bar{J}_{k}(\phi_{k}(i),r_{k})$ to the cost-to-go of image $\phi_{k}(i)$ at time period $k$ by adjusting the weights $r_{k}$ to minimize the sum of squares.
2.1 A Scalable Benchmark for Efficient Image Representations

### v2

# Optimal Control with Natural Images: Efficient Reinforcement Learning using Overcomplete Sparse Codes

@@ Abstract
Optimal control and sequential decision making are widely used in many complex tasks. Optimal control over a sequence of natural images is a first step towards understanding the role of vision in control. Here, we formalize this problem as a reinforcement learning task, and derive general conditions under which an image includes enough information to implement an optimal policy. Reinforcement learning is shown to provide a computationally efficient method for finding optimal policies when natural images are encoded into “efficient” image representations. This is demonstrated by introducing a new reinforcement learning benchmark that easily scales to large numbers of states and long horizons. In particular, by representing each image as an overcomplete sparse code, we are able to efficiently solve an optimal control task that is orders of magnitude larger than those tasks solvable using complete codes. Theoretical justification for this behaviour is provided. This work also demonstrates that deep learning is not necessary for efficient optimal control with natural images.
Keywords Natural image sequences $\cdot$ Optimal policies $\cdot$ Sufficient statistics $\cdot$ Sparse coding $\cdot$ Benchmark
@@ 1 Introduction
Many interesting and complex tasks can be described using natural image sequences. Each image of the sequence corresponds to a particular state of the environment, and choosing an action or applying a control at the current state takes the environment to its next state, and leads to the next image in the sequence. A natural question is then: given an image, what is the best control to apply? Although the answer will depend on the particular task at hand a general framework can be identified. As a simple example, consider moving through a rainforest and taking images of flowers in order to locate a particular species of plant. In this case, each image describes a state of the environment, and the state changes depending on the location of the next image. Each image also has a cost or reward associated with it: an image of the sought-after flower will receive a high reward, while an image without any flowers may receive a low (or even negative) reward. To perform this task well, the choice of control leading to the location of the next image does not necessarily give the best chance of obtaining an image of the sought-after flower at the next time period, rather, it should also lead to a good chance of obtaining such images at future time periods as you move through the rainforest. Planning over a sequence of time periods is the defining characteristic of sequential decision making and optimal control, and is a clear requirement of any intelligent system.
@@
Natural images form a high-dimensional dataset with a great deal of expressive power. For some tasks, images may include enough task-dependent information to determine an optimal policy. In this case no additional information is required for optimal control. In other cases, this might not be true and additional information may be necessary. Here, we will derive the general set of conditions for an image to be sufficient for implementing an optimal policy. Similar to other types of ecological data, it is well-known that the statistical properties of natural images distinguish them from artificially generated images (Field1; Field2; ruderman1; ruderman2; simoncellio; hyvarinenbook; Hosseini). We exploit these statistical properties to find efficient image representations called sparse codes. Sparse codes will then be shown to significantly reduce the computational resources required for finding optimal policies.
The aim of this work is to investigate the possibility of efficient optimal control over a sequence of natural images. The investigation has three stages. The first stage identifies what information an image must include so that optimal control is possible when an image of the environment is the only information available; in other words, when does an image become a “sufficient statistic”? The second stage identifies how to achieve efficient optimal control. Statistical dependencies in natural images lead to redundancy, making it possible to obtain an efficient representation using an appropriate encoding scheme such as sparse coding (daugman3; daugman2; olshausenfield1; olshausenfield2). A sparse code is “efficient” in the sense of being a low bit-rate (low entropy) representation of a high-dimensional dataset. Here, we show an overcomplete sparse code is also computationally efficient for finding optimal policies in reinforcement learning. The basic idea is described by the neural network shown in Figure 1. The first two layers of the neural network form a sparse autoencoder that generates an overcomplete sparse code by reconstructing an image input using an overcomplete basis of Gabor functions adapted to natural image statistics (loxley1). It is well-known that Gabor functions can be used to provide a “sparse description” of natural images (daugman3; daugman2). A linear regression network then makes use of this overcomplete sparse code to approximate the cost-to-go required in reinforcement learning. The efficiency gain is twofold: an overcomplete sparse code is shown to increase both the learning rate and the storage capacity of a linear network used for approximating cost-to-go values. The underlying theoretical reason is that a sparse code favourably conditions the Hessian matrix of the least squares problem, while an overcomplete sparse code also increases the rank of the associated design matrix. The third stage of the investigation empirically proves that an overcomplete sparse code gives the most efficient image representation out of several likely alternatives. This requires developing a new reinforcement learning benchmark to compare the performance and efficiency of different image representations. The benchmark must be scalable so the number of states and time periods can easily be increased; and optimal solutions must easily be distinguishable from suboptimal solutions. Scalability of the benchmark is necessary for determining which image representations will perform best over long horizons and for large numbers of states.
[…]
Figure 1: Neural network for reinforcement learning with natural images. The first two layers form a sparse autoencoder that generates an overcomplete sparse code $\phi$ by reconstructing the image input $I$ using an overcomplete basis of Gabor functions adapted to natural image statistics. The output $r\boldsymbol{\cdot}\phi$ then approximates the cost-to-go $\beta$ using weights $r$. The network storage capacity has increased from $d$ to (close to) $m$ by using an overcomplete sparse code.
@@
Previous work on reinforcement learning is heavily concentrated on synthetic datasets of limited complexity (although the tasks themselves may have very high complexity). Researchers have previously suggested that more work should be done to extend the paradigm of reinforcement learning to real-world data such as natural images and natural video (zhang2018natural). The few available examples of these efforts include using image representations to integrate unsupervised training of deep neural networks into reinforcement learning; specifically for visual navigation (lange), and simulated autonomous driving tasks (zhang2021learning). A little more work exists on the role of sparse codes and sparse representations in reinforcement learning, such as work to obtain better policies with neural networks on certain standard classic control tasks (Martha1; Martha2; Rafati). This line of work dates back to early research on coarse coding in reinforcement learning. One the other hand, none of these works considered sparse codes of images or other high-dimensional datasets of high complexity. The closest previous work appears in loxley2, where sparse representations of images were applied to find optimal policies for a specific control task. Unfortunately, this work was limited by the size of the task; which required a hand-annotated video dataset, and the type of task; which was restricted to deterministic controllable dynamics rather than a general Markov decision process. These limitations did not allow for results to be extended to large numbers of states and long horizons. In this work, a more general approach to optimal control is taken, and the development of a new reinforcement learning benchmark now allows results to be easily scaled up.
@@
The structure of this paper is as follows. In Sec. 2, reinforcement learning methodology is briefly summarized, and the problem of optimal control over a sequence of natural images is formalized. Following this, a new reinforcement learning benchmark is developed in Sec. 3. In Sec. 4, results for the reinforcement learning benchmark are presented, and alternative image representations are compared. A detailed discussion of results and conclusions is given in Sec. 5.
@@ 2 Reinforcement Learning over Natural Image Sequences
We consider a general optimal control task involving a discrete time dynamical system with a finite number of states and controls. During any time period $k$, the environment is described by a state $i$ taken from a state space $S$, and a control $u$ is selected from a set of available controls $U_{k}(i)\subset C$ (we adopt the notation from bert). The state of the environment changes when a controller or agent chooses control $u$ from state $i$, moving the environment to state $j$ according to the transition probabilities $p_{ij}(u)$ of a controllable Markov chain. Since the controller or agent can choose different controls depending on the current state of the environment, the controller’s choice is described by a policy $\pi=\{\mu_{0},...,\mu_{N-1}\}$, where the sequence of functions $\mu_{k}$ map states into controls: $\mu_{k}(i)=u$, for each state $i$ and each time period $k$. The cost of being in state $i_{k}$ during time period $k$, choosing control $u_{k}$, then moving to state $i_{k+1}$ during time period $k+1$, is denoted by $g_{k}(i_{k},u_{k},i_{k+1})$. The key quantity of interest is the expected total cost. Given a policy $\pi$ and an initial state $i_{0}$, the expected total cost is the expected value of the sum of costs:
@@
The finite-horizon optimal control task outlined in Eqs. (1) and (2) can be solved exactly using the Dynamic Programming Algorithm, leading to the optimal cost-to-go $J_{k}^{*}$ (this algorithm is called Value Iteration in the case of infinite horizon tasks, and converges to $J^{*}$ for discounted problems with stationary policies). The optimal policy $\mu_{k}^{*}(i)$ is then given by the minimizing control $u^{*}$ in the expression:
@@
for each state $i$ and each time period $k$.
Optimal control with natural image sequences requires a slight modification of this general framework. The key consideration is now what information to include in a state to allow for optimal decision making. Any state that includes all necessary information for optimal control is called a sufficient statistic (bert). The assumption we make here is that an image is a sufficient statistic. In this case, each state $i\in S$ in an optimal control task is now given by an image $\phi(i)\in\bar{S}$; where $\bar{S}\subset\mathbb{R}^{p}$ so that each image is represented as a $p$-dimensional vector. The index $i$ used to denote a particular state in the general framework now denotes a particular image (state) in the modified framework. However, this index is simply a notational convenience to distinguish between different images, and is not part of the information required for optimal control. An image $\phi(i)$ is a sufficient statistic if $U_{k}(i),p_{ij}(u)$, and $g_{k}(i,u,j)$ in Eq. (3) depend on $i$ or $j$ only through $\phi(i)$ or $\phi(j)$, respectively. For example, a striking consequence of $\phi(i)$ as a sufficient statistic is the form the transition probabilities now take:
$p_{ij}(u)=\bar{p}(\phi(j)|\phi(i),u),$ (4)
so that given image $\phi(i)$, selecting control $u$ leads to image $\phi(j)$ with probability $\bar{p}(\phi(j)|\phi(i),u)$. Therefore, a controllable Markov chain now becomes an “image generator” when used to generate new samples. We derive a simple “image generator” in the next section. The main consequence of $\phi(i)$ as a sufficient statistic is that the optimal policy can now be written as
$\mu_{k}^{*}(i)=\bar{\mu}_{k}^{*}(\phi(i)),$ (5)
and so optimal control depends only on the information contained in the image $\phi(i)$.
Now that the general framework for optimal control has been appropriately modified to include natural image sequences, we present an algorithm for finding an optimal policy. Given an image $\phi(i)$ we require a function $\bar{J}_{k}(\phi(i))$, analogous to $J_{k}^{*}(i)$, that determines its cost-to-go. A simple choice of function is a linear network given by the scalar product of a weight vector $r_{k}\in\mathbb{R}^{p}$ and an image vector $\phi(i)$:
$\bar{J}_{k}(\phi(i),r_{k})=r^{\top}_{k}\phi(i),$ (6)
@@
where $\top$ is the vector transpose. To approximate the optimal cost-to-go $J_{k}^{*}$ requires adjusting the network weights using a training rule and a dataset. Here, we apply the algorithm known as Fitted Value Iteration. To start the iteration, we go to the end of the task at k=N1k=N-1, and make use of the known terminal cost $\bar{J}_{N}(\phi(i),r_{N}^{*})=g_{N}(i)$. The following pair of expressions are then iterated backwards from k=N1k=N-1 to $k=0$ using $\bar{J}_{N}$ to start the iteration,
$\beta_{k}^{i}=\underset{u\in\bar{U}_{k}(\phi(i))}{\operatorname{min}}\ \sum_{j}\bar{p}(\phi(j)|\phi(i),u)\left\{\bar{g}_{k}(\phi(i),u,\phi(j))+\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})\right\},\ \ \ \ \ \forall i$ (7)
$\text{minimize}\ \ \ \sum_{i}(\bar{J}_{k}(\phi(i),r_{k})-\beta_{k}^{i})^{2},$ (8)
$\text{subject to}\ \ \ r_{k}\in\mathbb{R}^{p},$
@@
yielding the cost-to-goes $\bar{J}_{N-1},...,\bar{J}_{0}$ required for finding an optimal policy. The first expression (7) yields the cost-to-go $\beta_{k}^{i}$ of image $\phi(i)$ at time period $k$ by making use of the cost-to-go $\bar{J}_{k+1}(\phi(j),r_{k+1}^{*})$ of image $\phi(j)$ at time period $k+1$, as well as the current cost $\bar{g}_{k}(\phi(i),u,\phi(j))$. In the second expression (8), the “data” given by $(\phi(i),\beta_{k}^{i})$ from Eq. (7) is used to fit the linear network $\bar{J}_{k}(\phi(i),r_{k})$ to the cost-to-go $\beta_{k}^{i}$ by adjusting the weights $r_{k}$ to minimize the sum of squares.
Natural images used for optimal control in (6)–(8) are computationally inefficient when using raw image pixels, as will be demonstrated in the next section. It turns out to be much better to use some type of image representation. The sparse autoencoder in Figure 1 generates one possible representation for natural images, where $\phi(i)\in\mathbb{R}^{p}$ is given by an overcomplete sparse code (i.e., $p=m>d$). In this case, training the linear network for optimal control requires solving two linear least-squares problems: 1) minimize $\lVert G\phi-I\rVert_{2}^{2}$ to find an overcomplete sparse code $\phi\in\mathbb{R}^{m}$ given an image input $I\in\mathbb{R}^{d}$ and an overcomplete basis of Gabor functions $G\in\mathbb{R}^{d\times m}$ (adapted to natural image statistics), followed by 2) minimize $\lVert r\boldsymbol{\cdot}\phi-\beta\rVert_{2}^{2}$ to find a weight vector $r\in\mathbb{R}^{m}$ after making use of (7) to find $\beta$. This must be done for each image input.
3 A Scalable Benchmark for Efficient Image Representations